OpenStudy (anonymous):

Solve y over negative 2+ 5 = −4

OpenStudy (anonymous):

2 −2 18 −18

OpenStudy (anonymous):

y =−18

OpenStudy (anonymous):

got time for a couple more

OpenStudy (anonymous):

In a certain county, the number of Charter Schools is ten less than twice the number of Alternative Schools. We know that there are 46 Charter Schools in the county. How many Alternative Schools are in the county?

OpenStudy (texaschic101):

c = charter schools a = alternative schools c = 2a - 10 c = 46 46 = 2a - 10 46 + 10 = 2a 56 = 2a 56/2 = a 28 = a check c = 2a - 10 46 = 2(28) - 10 46 = 56 - 10 46 = 46 (correct) There are 28 alternative schools

OpenStudy (anonymous):

thanks you

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

can you answer another

OpenStudy (texaschic101):

possibly...what you got

OpenStudy (anonymous):

Solve: 3(x − 1) − (x + 5) = 2

OpenStudy (texaschic101):

3(x - 1) - (x + 5) = 2 3x - 3 - x - 5 = 2 2x - 8 = 2 2x = 2 + 8 2x = 10 x = 5

OpenStudy (texaschic101):

I just used the distributive property

OpenStudy (anonymous):

i have one more The sum of two consecutive even integers is 234. What is the larger integer?

OpenStudy (texaschic101):

x is one number x + 2 is the 2nd number x + x + 2 = 234 2x + 2 = 234 2x = 234 - 2 2x = 232 x = 232/2 x = 116 x + 2 = 118 <--larger number

OpenStudy (anonymous):

ur good at this

OpenStudy (texaschic101):

I try :)

OpenStudy (anonymous):

thats all im going to ask

OpenStudy (texaschic101):

ok....glad I could help :)

OpenStudy (anonymous):

:)